# coverforest: Conformal Predictions with Random Forest in Python

### v1

@@ 3.2 Source code
The source code for the coverforest is hosted on GitHub at https://github.com/donlapark/coverforest. The code follows the PEP8 guidelines, and all docstrings of publicly exposed classes comply with the NumPy docstring style. Both CoverForestClassifier and CoverForestRegressor pass the comprehensive unit tests provided by scikit-learn. The documentation includes installation guides, user manual, and API reference, generated using sphinx and available online at https://donlapark.github.io/coverforest. The package is BSD-licensed and requires only scikit-learn version 1.6.1 or greater.
@@ 4 Examples
We train a conformal random forest using CV+ with the RAPS score. To use the CV+ method, initialize a CoverForestClassifier model with method="cv". Here, we set cv=5 to split the data into five folds. As a result, five random forests will be fitted on the five combinations of 4-folds. We also specify that each forest has 200 trees by setting n_estimators=200. {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=]python from coverforest import CoverForestClassifier from coverforest.metrics import average_set_size_loss from coverforest.metrics import classification_coverage_score from sklearn.datasets import load_digits from sklearn.model_selection import train_test_split
X, y = load_digits(return_X_y=True) X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2)
@@
We train the random forest by calling the fit() method, and make a prediction by calling the predict() method. During prediction, we choose the miscoverage rate at $0.05$ by specifying alpha=0.05. The output of the predict() method is a tuple: The first item contains the single class predictions by the random forest, and the second item contains the set predictions. {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=]python clf.fit(X_train, y_train) y_pred, y_sets = clf.predict(X_test, alpha=0.05) # y_pred[:5] = [2, 8, 2, 6, 6] # y_sets[:5] = [array([2]), array([2, 5, 8]), array([2]), # array([5, 6]), array([6])]
@@
We may use the provided metrics functions to evaluate the set predictions against the true classes. {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=]python avg_set_size = average_set_size_loss(y_test, y_sets) coverage = classification_coverage_score(y_test, y_sets) # avg_set_size = 1.589 # coverage = 0.964 The small average set size indicates that the model can make predictions with high confidence. The coverage tells us that $96.4\%$ of the prediction sets contain the true classes, which is above our specified level of $95\%$.
@@
In the second example, we train a conformal random forest using Jackknife+-after-bootstrap (J+ab) with the APS score. To use J+ab, initialize a CoverForestClassifier model with method="bootstrap". To use the APS method, set k_init=0 and lambda_init=0. In contrast to CV+ which has multiple forests, J+ab has only a single forest. Here, we specify that this forest has 400 trees. {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=]python clf = CoverForestClassifier( n_estimators=400, method="bootstrap", k_init=0, lambda_init=0 ) clf.fit(X_train, y_train) y_pred, y_sets = clf.predict(X_test, alpha=0.05) avg_set_size = average_set_size_loss(y_test, y_sets) coverage = classification_coverage_score(y_test, y_sets) # y_pred[:5] = [2, 8, 2, 6, 6] # y_sets[:5] = [array([2]), array([2, 5, 8, 9]), # array([2]), array([6]), array([6])] # avg_set_size = 1.761 # coverage = 0.968 Compared to the CV+ method, J+ab produces prediction sets with a larger average with practically the same coverage.
@@ 5 Experiments and Discussions
To validate our implementation, we benchmark coverforest’s three methods: CV+, Jackknife+-after-bootstrap (J+ab), and split conformal on conformal classfication and regression tasks.
@@ 5.1 Classification experiment
We evaluate the predictions and run times of CoverForestClassifier on four classification datasets. See Table 2 for the datasets’ characteristics.
To evaluate a model on each dataset, we split the data with 80% as the training set and 20% as the test set; then we fit the models on the training set with the target miscoverage level $\alpha\in\{0.05,0.1,0.2\}$. We repeat the experiment 50 times for each method and compare the empirical coverage (the proportion of prediction sets that contain the true class), and average set size. For fair comparisons, we balance the number of trees between the three methods so that they have the same predictive power. This involves ensuring that each training sample is be used to train the same number of trees across all these models. For CV+, the data is split into 10 folds, and we fit 10 forests with 100 trees each, totaling 1000 trees. Since each tree in CV+ is trained on 90% of the data, we accordingly set both J+ab and split conformal methods to use 900 trees to maintain consistency in the learning capacity.
Instances Classes Features Training set Test set
Mice [26] 1080 8 80 864 216
WineQuality [27] 2000 7 11 1600 400
Myocardial [28] 1700 8 111 1360 240
MNIST [29] 1250 10 784 1000 250
@@
Table 2: The datasets used in the conformal classification experiment.
@@
Figure 2: Coverage probabilities of CoverForestClassifier’s set predictions on the test set, repeated 50 times for each of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the APS and RAPS scores.
Empirical coverage
Figure 2 shows the empirical coverages for all methods using both APS (4) and RAPS (5) scores. In all datasets, the median coverages generally align with the nominal levels. What is most surprising is the MNIST datasets, where all methods achieve the coverages way above the nominal levels.
In the Mice and Myocardial datasets, there is no difference between the APS and RAPS scores. However, in the WineQuality and MNIST datasets, methods with the APS score typically achieve higher coverage than those with the RAPS score, with CV+ methods consistently outperforming J+ab in terms of coverage.
Average set size
Figure 3 shows the average set sizes across the methods. In all datasets, split conformal exhibits notably higher variance in set sizes compared to CV+ and J+ab methods. The RAPS score’s regularization effectively reduced set sizes in the Mice, Myocardial, and MNIST datasets, with the most pronounced reduction observed in MNIST. This substantial reduction in MNIST can be attributed to the visual similarity between certain digit pairs (e.g., 4/9, 3/8, 1/7), which causes the model to distribute non-negligible probabilities across multiple classes even for clear instances. In both WineQuality and MNIST datasets, J+ab consistently produced smaller prediction sets than CV+, corresponding to its lower coverage rates observed in these datasets.
@@
Figure 3: Average size of CoverForestClassifier’s set predictions on the test set, repeated 50 times for each of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the APS and RAPS scores.
Run time comparison with MAPIE
In this experiment, we compare the run time of CoverForestClassifier against MAPIE [30], a Python package that implement wrapper classes that enable conformal prediction on any scikit-learn models. It is currently the only package that provides the CV+ and J+ab methods with the APS and RAPS scores (though J+ab with APS and RAPS for conformal classification has not been implemented). MAPIE’s development emphasizes versatility across all machine learning models, whereas coverforest’s development focuses on efficient random forest implementations.
@@
For coverforest and MAPIE, we train a random forest and make conformal predictions using CV+, J+ab and split conformal methods, all with the APS score. We fit the models across the four datasets introduced in Experiment 1. We now use the full WineQuality dataset of 3918 training and 980 test samples, and the MNIST dataset is expanded to 5000 training and 1250 test samples. We run the three methods with the same number of trees as in Experiment 1 (1000 trees for CV+ and 900 trees for J+ab and split conformal). And we fix $\alpha=0.05$. We run each method for 30 repetitions each and report the average training and prediction times.
Figure 4: Training and prediction times of coverforest’s and MAPIE’s implementations of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the APS score and $\alpha=0.05$.
@@
Figure 4 compares the run times of both packages (note that at the time of this study, MAPIE had not implemented J+ab with the APS score). It shows that coverforest consistently runs faster than MAPIE in both training and prediction, especially with large datasets like WineQuality and MNIST. For the CV+ method specifically, coverforest’s optimized implementation reduces cross-conformal prediction times to only a few seconds.
@@ 5.2 Regression experiment
Instances Output Range Features Training set Test set
Housing [31] 5000 [0.15, 5.0] 8 4000 1000
Concrete [32] 1030 [2.33, 82.6] 8 824 206
Bike [33] 5000 [1, 638] 12 4000 1000
Crime [34] 1994 [0.0, 1.0] 122 1595 399
@@
We use the same numbers of trees as in Experiment 1 (1000 trees for CV+ and 900 trees for J+ab and split conformal). For each dataset, we run each method for 50 repetitions and calculate the coverage and average interval length. For benchmarking run times, we use the full Housing dataset (16,512 training, 4,128 test samples) and Bike dataset (13,903 training, 3,476 test samples). We fix $\alpha=0.05$. For each method, we perform training and prediction for 30 repetitions and compare the average run times to those of MAPIE.
Empirical coverage and average interval length
Figure 5: Coverage and average length of CoverForestRegressor’s interval predictions on the test set, repeated 50 times for each of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the residuals as the conformity scores.
Figure 5 shows the coverages and average interval lengths of the three methods. Across all datasets, both CV+ and J+ab produce prediction intervals with small variances compared to the split conformal, with J+ab consistently producing intervals of smaller average lengths and smaller coverages than CV+.
Run time comparison with MAPIE
Figure 6 shows the run times of coverforestand MAPIE on conformal regression. While prediction times are similar for both packages, coverforest demonstrates faster training times than MAPIE for all methods and datasets. The run time differences are small on smaller datasets like Concrete due to computational overheads. Conversely, on larger datasets, coverforest’s CV+ and split conformal train substantially faster than MAPIE’s implementations.
Figure 6: Training and prediction times of coverforest’s and MAPIE’s implementations of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the residuals as the conformity scores and $\alpha=0.05$.

### v2

@@ 3.3 Source code
The source code for the coverforest is hosted on GitHub at https://github.com/donlap/coverforest. The code follows the PEP8 guidelines, and all docstrings of publicly exposed classes comply with the NumPy docstring style. Both CoverForestClassifier and CoverForestRegressor pass the comprehensive unit tests provided by scikit-learn. The documentation includes installation guides, user manual, and API reference, generated using sphinx and available online at https://donlapark.github.io/coverforest. The package is BSD-licensed and requires only scikit-learn version 1.6.1 or greater.
@@ 4 Examples
We train a conformal random forest using CV+ with the RAPS score. To use the CV+ method, initialize a CoverForestClassifier model with method="cv". Here, we set cv=5 to split the data into five folds. As a result, five random forests will be fitted on the five combinations of 4-folds. We also specify that each forest has 200 trees by setting n_estimators=200.
from coverforest.metrics import average_set_size_loss
from coverforest.metrics import classification_coverage_score
from sklearn.datasets import load_digits
from sklearn.model_selection import train_test_split
X, y = load_digits(return_X_y=True)
X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2)
@@
We train the random forest by calling the fit() method, and make a prediction by calling the predict() method. During prediction, we choose the miscoverage rate at $0.05$ by specifying alpha=0.05. The output of the predict() method is a tuple: The first item contains the single class predictions by the random forest, and the second item contains the set predictions.
y_pred, y_sets = clf.predict(X_test, alpha=0.05)
# y_pred[:5] = [2, 8, 2, 6, 6]
# y_sets[:5] = [array([2]), array([2, 5, 8]), array([2]),
# array([5, 6]), array([6])]
@@
We may use the provided metrics functions to evaluate the set predictions against the true classes.
coverage = classification_coverage_score(y_test, y_sets)
# avg_set_size = 1.589
# coverage = 0.964
The small average set size indicates that the model can make predictions with high confidence. The coverage tells us that $96.4\%$ of the prediction sets contain the true classes, which is above our specified level of $95\%$.
@@
In the second example, we train a conformal random forest using J+ab with the APS score. To use J+ab, initialize a CoverForestClassifier model with method="bootstrap". To use the APS method, set k_init=0 and lambda_init=0. In contrast to CV+ which has multiple forests, J+ab has only a single forest. Here, we specify that this forest has 400 trees.
n_estimators=400, method="bootstrap", k_init=0, lambda_init=0
clf.fit(X_train, y_train)
y_pred, y_sets = clf.predict(X_test, alpha=0.05)
avg_set_size = average_set_size_loss(y_test, y_sets)
coverage = classification_coverage_score(y_test, y_sets)
# y_pred[:5] = [2, 8, 2, 6, 6]
# y_sets[:5] = [array([2]), array([2, 5, 8, 9]),
# array([2]), array([6]), array([6])]
# avg_set_size = 1.761
# coverage = 0.968
Compared to the CV+ method, J+ab produces prediction sets with a larger average with practically the same coverage.
@@ 5 Experiments and Discussions
To validate our implementation, we benchmark coverforest’s three methods: CV+, J+ab, and split conformal on conformal classfication and regression tasks. We compare our implementations against the most recent packages for conformal predictions, namely the crepes and MAPIE packages.
@@ 5.1 Classification experiment
We evaluate the predictions and run times of CoverForestClassifier on five classification datasets, namely Mice [32], WineQuality [33], Bean [34], MNIST, and Helena [35] datasets. See Table 3 for the datasets’ characteristics.
We aim to evaluate the methods’ abilities to achieve the desired level of coverage even with a few training points. To this end, we randomly select 20% of the dataset as the test set (except the MNIST dataset whose test set has already been prepared), and 200 rows of the remaining data as the training set. For the split conformal method, we reserve 100 rows of the training set as the calibration set. We then fit the models on the training set with the target miscoverage level $\alpha\in\{0.05,0.1,0.2\}$. We repeat the experiment 50 times for each method and compare the empirical coverage (the proportion of prediction sets that contain the true class), and average set size.
For CV+, we split the training data into 10 folds, and each random forest has 30 trees. For J+ab and split conformal, the random forest has 100 trees.
Experiment 1 & 2 Experiment 3
Classes Features Train Test Train Test
Mice [32] 8 80 200 216 861 216
WineQuality [33] 7 11 200 400 3918 980
Bean [34] 7 16 200 2723 5000 1250
MNIST [36] 10 784 200 10000 5000 1250
Helena [35] 100 27 1033 300 904 300
@@
Table 3: The datasets used in the conformal classification experiments. The smaller training set of the Helena dataset in Experiment 3 is due to the memory limit of the MAPIE package.
We also compare our methods against crepes [37], which uses the classifier’s predicted probabilities as the conformity scores; however, such scores are not adaptive in the sense that they tend to result in coverages that are over or under the nominal level, depending on the accuracy of the classifier.
crepes offers the OOB and split conformal methods. In our experiment, we run both methods with a random forest with 100 trees as the base model.
As the MAPIE library implements the same CV+ and split conformal methods as ours, we do not run MAPIE in this experiment. Nonetheless, we will compare our run times against MAPIE in a later experiment.
For Experiment 1 and 2, we also perform hypothesis testing to verify that the empirical coverage probabilities exceed the nominal coverage levels, and to compare the average set sizes between any pair of the methods. The results and discussion of the tests can be found in A.
@@
Figure 2: Coverage probabilities of CoverForestClassifier’s set predictions on the test set, repeated 50 times for each of the CV+, J+ab and split conformal methods with the APS and RAPS scores. The results of crepes’s methods are also included.
Experiment 1. Empirical coverage
Figure 2 presents the empirical coverages achieved using both the APS (4) and RAPS (7) scores. The plots consistently demonstrate that CV+ provides the largest coverage. J+ab consistently yields slightly lower coverages than CV+. On the Mice, Bean, and Helena datasets, the OOB method from crepes results in notably large coverage across all values of $\alpha$. In contrast, the split conformal method provides the lowest and most inconsistent coverages on these same datasets.
We now turn our attention to the effect of regularization on the empirical coverages. In this regard, the CV+ and J+ab methods, when paired with the RAPS score, yield slightly reduced coverages for $\alpha=0.1$ and $0.2$. This difference becomes more pronounced in datasets with a large number of classes, as is evident in the results for the MNIST and Helena datasets.
Finally, we assess whether these methods achieve their nominal coverage levels. The plots indicate that the median coverages for both the CV+ and J+ab methods exceed these target levels in the majority of cases. This outcome demonstrates the data efficiency of these two approaches, which allows them to successfully reach the desired coverage.
Experiment 2. Average set size
Figure 3 illustrates the average prediction set sizes produced by each method. Across all datasets, we observe a general trend where CV+ produces the largest average set sizes, followed by J+ab, and then the split conformal method. The split and OOB methods from crepes yield the smallest set sizes on the Mice, Bean, and MNIST datasets. It is important to recall from the previous experiment, however, that these methods did not achieve the nominal coverage levels on the MNIST dataset.
We now examine the impact of regularization on the average set sizes. For all datasets, the application of the RAPS score contributes to a reduction in the set sizes for the CV+, J+ab, and split conformal methods. This decrease in set size is substantially more noticeable for datasets with a large number of classes, as demonstrated by the results on the Helena dataset.
@@
Figure 3: Average size of CoverForestClassifier’s set predictions on the test set, repeated 50 times for each of the CV+, J+ab and split conformal methods with the APS and RAPS scores. The results of crepes’s methods are also included.
Experiment 3. Run time comparison with MAPIE
In this experiment, we compare the run time of CoverForestClassifier against MAPIE [31]. As far as we know, it is currently the only package that provides the CV+ methods with the APS and RAPS scores (the J+ab method is also provided in MAPIE but only for regression tasks).
We do not compare our methods against the crepes library here as it does not provide the CV+ and J+ab methods, nor the APS and the RAPS scores.
@@
For coverforest and MAPIE, we train a random forest and make conformal predictions using CV+, J+ab and split conformal methods, all with the APS score. The numbers of training and test instances in this experiment are indicated in Table 3. We run CV+ with 100 forests with 10 trees each, and J+ab and split conformal with 900 trees. And we fix $\alpha=0.05$. We run each method for 30 repetitions each and report the average training and prediction times.
@@
Figure 4 compares the run times of both packages (note that at the time of this study, MAPIE had not implemented J+ab with the APS score). It shows that coverforest consistently runs faster than MAPIE in both training and prediction, especially with large datasets like WineQuality and MNIST. The improvements are even more pronounced on datasets with many classes. Specifically, on the Helena dataset, coverforest training time is 14 times faster for the CV+ method, 8 times faster for the split method; and the testing time is 72 times faster for the CV+ method, and 30 times faster for the split method. These results indicate the performance advantage to our optimized Numpy code during the training phase, and the optimized Cython code during the prediction phase.
Figure 4: Training and prediction times of coverforest’s and MAPIE’s implementations of the CV+, J+ab and split conformal methods with the APS score and $\alpha=0.05$. Note that MAPIE does not implement the J+ab method for classification.
@@ 5.2 Regression experiment
Experiment 1 & 2 Experiment 3
Ranges Features Train Test Train Test
Housing [38] [0.15, 5.0] 8 200 1000 16512 4128
Concrete [39] [2.33, 82.6] 8 200 206 824 206
Bike [40] [1, 638] 12 200 1000 13903 3476
Crime [41] [0.0, 1.0] 122 200 399 1595 399
@@
For CV+, we split the training data into 10 folds, and each random forest has 30 trees. For J+ab and split conformal, the random forest has 100 trees. We randomly sample 20% of each dataset for the test set and 200 instances of the remaining data for the training set. For each dataset, we run each method for 50 repetitions and calculate the coverage and average interval length. The numbers of training and test instances in this experiment are indicated in Table 4. We fix $\alpha=0.05$. For each method, we perform training and prediction for 30 repetitions and compare the average run times to those of crepes and MAPIE.
For Experiment 1 and 2, we also perform hypothesis testing to verify that the empirical coverage probabilities exceed the nominal coverage levels, and to compare the interval lengths between any pair of the methods. The results and discussion of the tests can be found in A.
Experiment 1 & 2. Empirical coverage and average interval length
Figure 5: Coverage and average length of CoverForestRegressor’s interval predictions on the test set, repeated 50 times for each of the CV+, J+ab and split conformal methods with the residuals as the conformity scores. Results of crepes’s J+ab method are also included.
Figure 5 shows the coverages and average interval lengths of the three methods. Across all datasets, both CV+ and J+ab produce prediction intervals with small variances compared to the split conformal, with J+ab consistently producing intervals of smaller average lengths and smaller coverages than CV+. Even though both crepes’s methods provide coverages that achieve the nominal levels, their interval lengths are substantially longer than those from the other methods.
Experiment 3. Run time comparison with crepes and MAPIE
Figure 6 presents a comparison of the run times for coverforest, crepes, and MAPIE on conformal regression tasks. The results clearly show that coverforest is substantially faster than the other packages during the training phase across all evaluated methods.
Regarding prediction time, crepes is generally the fastest package, noticeably with the bootstrap (J+ab) method; this is due to the implementation of the Mondrian conformal method, which only requires a single prediction for each test point. In contrast, the methods implemented in our package and in MAPIE necessitate making multiple predictions for each test point (more precisely, for each training point, its out-of-bag trees are used to make predictions for the test point). Nonetheless, it is important to note that, unlike in classification, prediction in conformal regression takes substantially less time than training.
Figure 6: Training and prediction times of coverforest, crepes and MAPIE’s implementations of the CV+, J+ab and split conformal methods with the residuals as the conformity scores and $\alpha=0.05$. Note that crepes does not implement the CV+ method.